Word Problems Workshee You have a jar of loose change which contains pen total value of the change in the jar is $104. In total, there are exactly 1000 coins in the jar There are four times as many pennies as there are nickels and there are twice as many pennies as there are dimes. How many of each type of coin are there? 4. nies, nickels, dimes, and quarters. The
Answer #1

Suppose p, n, d and q are denoting number of pennies, nickels, dimes and quarters present in the jar

Now there are 1000 coins in the jar. p+n+d+q-1000......(1)

The total value of change in the jar is $104 0.01p + 0.05n + 0.10d + 0.25q = 104   ......(2)

There are four times as many pennies as there are nickels
n=0.25p ......(3)

There are twice times as many pennies as there are dimes
d = 0.5p ......(4)


Now,,values,,of,,n,,and,,d,,from,,the,,equations,,(3),,and,,(4),,intoequations,,(1),,and,,(2)
p+ 0.25p +0.5p q 1000 1.75p q 1000 q 1000 1.75p ......(5)

0.01p+0.05 imes 0.25p+0.10 imes0.5p+0.25q=104 Rightarrow ,,0.0725p+0.25q=104 Rightarrow ,,0.0725p+0.25 imes(1000-1.75p)=104q = 1000-1.75p from eq.(5)
Rightarrow ,,0.0725p+250-0.4375p=104 Rightarrow ,,0.4375p-.0725p=250-104 Rightarrow ,,0.365p=146 Rightarrow ,,p=rac{146}{0.365} Rightarrow p=400 ......(6)

From,,eq.(6),,and,,(3) n=0.25p Rightarrow ,,n=0.25 imes400 Rightarrow ,,n=100

From eq. (6) and (4) d = 0.5p d = 0.5 × 400 d=200

From,,eq.(6),,and,,(5) q=1000-1.75p Rightarrow ,,q=1000-1.75 imes400 Rightarrow ,,q=300

Therefore,,,
  Number of pennies p = 400
     nickels n = 100
dimes,,d=200
quarters q = 300